OpenStudy (anonymous):

find y' and y''. y=(ln x)/(x^2)

OpenStudy (jack1):

if \[y = \frac {\ln x}{x^2}\] u can also write it as \[y = \ln x \times x^{-2}\] if it makes it easier for u to think about it? so are u familiar with the product rule? @muneka.fea2956 ? if u call ln x --> "u" and x^-2 --> "v" product rule states: \[y = v \ \times u \rightarrow so \rightarrow y' = (u' \times v) + (v' \times u)\] is this makin sense so far? sorry, it takes me a while to type in latex

OpenStudy (jack1):

so back to the above: v = x^-2 u = ln x v' = -2x^-3 u' = 1/x so add them all together as per the product rule y′=(u′×v)+(v′×u) = (1/x) times (1/x^2) + ( -2/x^3) times (ln x) = 1/x^3 + -2 ln x / x^3 = 1 - 2 ln x -------------- x^3 or eqn form: \[y' = \frac {1 - 2 \ln x}{x^3} \]

OpenStudy (jack1):

so y'' ... just use product rule again to find derivative of above @muneka.fea2956

OpenStudy (jack1):

this time u = (1 - 2 ln x) and v = x^-3

OpenStudy (anonymous):

thank you!! :)

OpenStudy (jack1):

all good ;D